# SpikingBrain Technical Report: Spiking Brain-inspired Large Models

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) built on the Transformer architecture (vaswani2017attention, ) have been driven by the scaling law (kaplan2020scaling, ), which suggests that performance improves with larger model sizes and more data (OpenAI2025GPT5, ; GoogleDeepMind2024GeminiPro, ; Anthropic2025ClaudeOpus41, ; liu2024deepseek, ). However, this scale-driven approach comes with significant challenges: extremely high training costs, high energy consumption, and complex deployment pipelines. Therefore, achieving high performance and energy efficiency under limited resources has become a critical research goal. To address this, our work draws inspiration from brain mechanisms. We explore novel architectures, training paradigms, and spike coding schemes to develop efficient, brain-inspired LLMs that move beyond the traditional Transformer framework.
A key focus of this study is to validate the training and deployment of such models on non-NVIDIA computing clusters. We use an open-source Transformer checkpoint (Qwen2.5-7B-base (yang2024qwen2, ) as an example) together with our efficient development framework to train and evaluate two models on the MetaX GPU cluster. The models, SpikingBrain-7B and SpikingBrain-76B-A12B, undergo end-to-end validation, including continual pre-training (CPT), long-context extension (up to 128k tokens), and supervised fine-tuning (SFT). We also adapt the vLLM inference framework to demonstrate deployment feasibility on MetaX hardware.
Our main technical contributions are as follows:
@@
In summary, this work presents efficient brain-inspired LLM training on MetaX GPUs, producing two models: SpikingBrain-7B and SpikingBrain-76B MoE (see Figure 1 for an overview). They achieve (near-)linear complexity, high training stability, and competitive performance with only $\sim 2\%$ of the pre-training data typical for mainstream LLMs. Inference shows substantial speedups, with SpikingBrain-7B reaching over 100× speedup in TTFT (Time to First Token) for 4M-token inputs.We further deploy a compressed 1B SpikingBrain model on a CPU-based mobile inference framework, achieving a 15.39× speedup at a sequence length of 256k. The proposed spiking scheme also yields $\sim 69\%$ sparsity, offering strong support for reducing power consumption. These findings highlight the potential of brain-inspired design for future neuromorphic hardware and efficient model scaling on diverse GPU platforms.
@@ 2.2 Integrated Model Architectures
Through lightweight training, a base Transformer can be converted into efficient attention variants of different forms (see Section 3). This enables flexible trade-offs between performance and efficiency by adjusting factors such as the proportion of full-attention layers or the window size of local attention. As a case study, we develop two models from the Qwen2.5-7B-base checkpoint: SpikingBrain-7B, a pure linear model optimized for long-context efficiency, and SpikingBrain-76B, a hybrid-linear MoE model designed to balance efficiency and performance. Their architectures are illustrated in Figure 2. Both models are integrated with HuggingFace and vLLM inference frameworks, supporting deployment on single- and multi-GPU settings.
@@
The 7B model achieves purely linear complexity by interleaving linear attention and sliding window attention (SWA) layers with a fixed 4K window in a 1:1 ratio. The FFN modules adopt the same SwiGLU design as the base model. In this architecture, SWA captures precise local patterns, while linear attention efficiently compresses long-range information. The model thus provides linear-time training complexity and constant memory usage during inference, regardless of sequence length, offering substantial efficiency gains for long-context processing. In practice, we implement linear attention using a Gated Linear Attention Module yang2024gated (GLA), where the gating vector $\mathbf{g}_{t}$, derived from low-rank projection parameters, enhances expressivity and recurrent modeling capability:
@@
The 76B model integrates linear attention and SWA in a 1:1 intra-layer parallel hybrid configuration, while standard full-attention layers are interleaved at a 1:6 ratio across layers. During parallel hybridization, outputs of both attention branches are normalized to ensure consistent scale and to prevent instability in the early stages of training. In addition, 128 learnable sink tokens (xiao2023efficient, ; dong2024hymba, ) are introduced to mitigate the attention-sink phenomenon in softmax attention and to enhance the flexibility of SWA for local modeling. Specifically, 128 trainable embeddings are prepended to the input; these tokens are attended by all other tokens and also attend to each other without causal masking. This functionality is implemented by customizing FlashAttention (daoflashattention, ) kernels. The model also employs Gated Linear Attention Modules, but in this case, the gating vector is tied directly to the key vector (chou2024metala, ; qinhgrn2, ), i.e., $\mathbf{g}_{t}=1-\mathbf{k}_{t}$, eliminating the need for extra gating parameters. The FFN modules adopt a sparse Mixture-of-Experts (MoE) design: each MoE layer consists of 16 routed experts (top-1 activated) and one shared expert, so that only about 15% of parameters (12B) are activated per token. To stabilize training and control parameter growth, 7 dense FFN layers are preserved out of 28 total layers (specifically at layers 1, 2, 3, 5, 7, 9, and 11).
@@
Our architectural choices are closely aligned with principles observed in biological brains. i) Linear attention modules exhibit modeling properties analogous to human memory, relying on compressed and continuously updated states (gu2025tradeoffs, ). At each step, they retrieve information only from the current memory, showing Markov-like behavior. From a biological perspective, their stateful temporal recurrence can be viewed as a simplified abstraction of dendritic dynamics with multi-branch morphology. ii) The Mixture-of-Experts (MoE) component embodies the principle of modular sparse activation and functional specialization, reminiscent of the distributed and specialized processing found in neural circuits (o2021and, ). iii) Our spike coding scheme (see Section 3.3) draws inspiration from event-driven and adaptively sparse neuronal activation in biological systems (maass1997networks, ). By combining network-level sparsity (MoE) with neuron-level spiking sparsity, our approach enables on-demand allocation of computation and provides a robust two-scale efficiency strategy. Collectively, these results suggest a promising pathway for designing large-model architectures that are both efficient and biologically plausible (10636118, ).
@@ 3 Training Paradigm
3.1 Generality: Attention Map Correspondence
@@
Leveraging this attention-map correspondence, we can initialize the QKV projection parameters directly from a pre-trained Transformer checkpoint (kasai2021finetuning, ). By reusing the learned QK similarity and performing lightweight training on a small amount of data, the attention can be adapted to these local or low-rank cases. Since SWA maintains local modeling precision and linear attention retains global interaction, combining them in a hybrid paradigm provides a closer approximation to the original attention map (dass2023vitality, ; chen2021scatterbrain, ; arora2024simple, ). This enables a smoother transition during conversion and accelerates loss convergence.
Finally, due to the generality of this attention-map formulation, we can construct diverse attention patterns—sparse, local, linear, or hybrid—and transfer them from any pre-trained softmax attention model.
@@
3.2 Efficiency: Conversion-based Training
Multi-stage Conversion Pipeline
The continual pre-training process comprises three stages which progressively extend the context window while adapting the model. In the first stage, our models are trained on 100B tokens with a sequence length of 8k, aiming to transfer attention patterns toward local or low-rank variants and ensure loss convergence. Subsequently, the second and third stages extend the sequence length to 32k and 128k, respectively, each trained with 20B to 30B tokens. The entire conversion process consumes about 150B tokens in total. Compared to the $\sim$10T tokens typically required for training from scratch, the CPT approach requires only about 2% of the data, enabling efficient adaptation under resource and budget constraints. All three stages use the Matrix (zhang2024mapneo, ) dataset, with long-context data generated through a simple per-domain packing strategy. The RoPE base remains consistent with the base model at 1M.
MoE Upcycling
@@ 3.3.2 Step 2: Spike Coding
$\mathbf{s}_{\text{INT}}$ $=\sum_{t=1}^{T}\mathbf{s}_{t},\quad\mathbf{y}=V_{\text{th}}(\mathbf{x})\cdot\sum_{t=1}^{T}(\mathbf{W}\mathbf{s}_{t}),$ (26)
@@
  • Binary Spike Coding {0,1}: This is the most basic event-driven spike coding method. Each time a unit spike is fired (with a value of $1$), it represents an activation of the neuron state, and the spike count is accumulated over continuous time steps. This coding scheme is intuitive and simple, with low computational overhead, making it suitable for scenarios with very low spike counts and effectively reducing system complexity. However, when representing large counts, this coding often requires many time steps to complete the representation. Additionally, the lack of directional optimization for spike firing results in a higher firing rate, further limiting system energy efficiency.
@@
  • Bitwise Spike Coding: Bitwise coding can be viewed as an event sequence encoding method, where integer count values are expanded bit by bit into spike events over time steps. Each time step corresponds to one binary bit of the count value, significantly compressing the time dimension overhead. As shown in Figure 3 (c), this mechanism supports three implementation forms to accommodate different symbolic representations and precision requirements: i) Pure bitwise encoding, suitable for positive integers, provides extremely high temporal compression in high-count scenarios. For instance, a count of 256 requires 256 consecutive time steps in binary coding, 128 in ternary coding, but only 8 steps in 8-bit bitwise encoding. ii) Bidirectional bitwise encoding uses $\pm 1$ to represent each bit, replacing negative part with 1-1. This halves the time step emission rate and reduces the required number of time steps by one (e.g., 7 steps for a count of 256), while maintaining equivalence in representation. iii) Two’s complement encoding incorporates sign information into the highest bit, supporting both positive and negative counts while retaining binary simplicity and biological plausibility. Compared to ternary coding, which still requires many time steps for higher bit counts, the bitwise encoding scheme significantly compresses time steps, reducing total spike emissions by up to 8× (for an 8-bit count). This effectively reduces overall spike communication overhead and computational load while maintaining precision, making it particularly suitable for high-precision, low-power, and time-sensitive neuromorphic computing tasks.
@@ 4.2 Operator Adaptation
  • JIT Compilation Optimization: During Triton’s just-in-time compilation stage, the Gated Linear Attention (GLA) (yang2024gated, ) operators used in SpikingBrain are reorganized at the code level. By refining instruction pipelining and register allocation, the operators achieve a balance between memory access latency and computational density. This stage directly leverages Triton’s hardware-agnostic general optimization strategies, enabling dynamic and efficient execution of compute kernels.
@@ 4.3 Parallel Topology
The memory demands of training large language models often exceed the capacity of a single GPU. To make training such models feasible, it is essential to employ efficient and scalable distributed training techniques alongside other memory reduction methods. These approaches distribute computational and storage loads while maintaining training efficiency. In this section, we present the parallelization strategies and training topologies employed to train large brain-inspired models on the MetaX cluster.
@@
Pipeline parallelism divides the model into different stages by layers. Each GPU is assigned a subset of layers, and intermediate activations are communicated between GPUs during the forward and backward passes using peer-to-peer (p2p) communication. We employ the 1F1B (narayanan2021efficientllmmg, ; harlap2018pipedream, ) scheduling algorithm for efficient pipeline execution. Furthermore, incorporating Mixture-of-Experts (MoE) layers interspersed with dense FFN layers helps mitigate the memory imbalance issue, where the first pipeline stage typically exhibits higher memory usage compared to other stages.
@@
In Mixture-of-Experts models, most parameters reside within the expert networks. Expert parallelism (rajbhandari2022deepspeed, ) addresses this by partitioning experts across multiple GPUs. This requires all-to-all communication to dispatch and combine tokens. We employ Grouped GEMM kernels to accelerate computation across experts, and an auxiliary loss is used to promote balanced token routing across experts.

### v2

# SpikingBrain: Spiking Brain-inspired Large Models

@@ 1 Introduction
Recent advances in large language models (LLMs) built on the Transformer architecture (vaswani2017attention, ) have been driven by the scaling law (kaplan2020scaling, ), which suggests that performance improves with larger model sizes and more data (OpenAI2025GPT5, ; GoogleDeepMind2024GeminiPro, ; Anthropic2025ClaudeOpus41, ; liu2024deepseek, ). However, this scale-driven approach comes with significant challenges: extremely high training costs, substantial energy consumption, and complex deployment pipelines. Therefore, achieving high performance and energy efficiency under limited resources has become a critical research goal. To address this, our work draws inspiration from brain mechanisms. We explore novel architectures, training paradigms, and spike coding schemes to develop efficient, brain-inspired LLMs that move beyond the traditional Transformer framework.
A further objective is to validate the training and deployment of such models on non-NVIDIA computing clusters. We use an open-source Transformer checkpoint (Qwen2.5-7B-base (yang2024qwen2, ) as an example) together with our efficient development framework to train and evaluate two models on the MetaX GPU cluster. The models, SpikingBrain-7B and SpikingBrain-76B-A12B, undergo end-to-end validation, including continual pre-training (CPT), long-context extension (up to 128k tokens), and supervised fine-tuning (SFT). We also adapt the vLLM inference framework to demonstrate deployment feasibility on MetaX hardware.
As illustrated in Figure 1, SpikingBrain constructs model architectures aligned with brain-inspired mechanisms; these architectures are developed via an efficient conversion pipeline and supported by adaptations for non-NVIDIA clusters. Our main technical contributions are as follows:
@@
In summary, this work presents efficient brain-inspired LLM development on MetaX GPUs, producing two models: SpikingBrain-7B and SpikingBrain-76B MoE. They achieve (near-)linear complexity, high training stability, and competitive performance with only $\sim 2\%$ of the pre-training data typical for mainstream LLMs. Inference shows substantial speedups, with SpikingBrain-7B reaching over 100× speedup in TTFT (Time to First Token) for 4M-token inputs. We further deploy a compressed 1B SpikingBrain model on a CPU-based mobile inference framework, achieving a 15.39× speedup at a sequence length of 256k. The proposed spiking scheme also yields $\sim 69\%$ sparsity, offering strong support for reducing power consumption. These findings highlight the potential of brain-inspired design for future neuromorphic hardware and enable efficient model scaling across diverse computing platforms (see Figure 2).
Figure 2: Compatibility of SpikingBrain models across diverse computing platforms. SpikingBrain models can be deployed on CPUs and both NVIDIA and non-NVIDIA GPUs using integer activation formats, and they also inspire the design of dedicated neuromorphic hardware leveraging event-driven sparse spike representations.
@@ 2.2 Integrated Model Architectures
Through lightweight training, a base Transformer can be converted into efficient attention variants of different forms, alongside MoE upcycling and spiking neuron coding (see Section 3). As a case study, we develop two models from the Qwen2.5-7B-base checkpoint: SpikingBrain-7B, a pure linear model optimized for long-context efficiency, and SpikingBrain-76B, a hybrid-linear MoE model designed to balance efficiency and performance. Their architectures are illustrated in Figure 3. Both models are integrated with HuggingFace and vLLM inference frameworks, supporting deployment on single- and multi-GPU settings. Additionally, as shown in Figure 2, SpikingBrain models can be deployed across diverse computing platforms, including CPUs as well as both NVIDIA and non-NVIDIA GPUs. Moreover, by expanding integer activations into equivalent sparse spike representations, the models provide insights that can inspire the design of next-generation neuromorphic chips.
@@
The 7B model achieves purely linear complexity by interleaving linear attention and sliding window attention (SWA) layers with a fixed 4K window in a 1:1 ratio. The FFN modules adopt the same SwiGLU design as the base model. Spike coding is applied to convert the activations of all linear projection layers in the model, as matrix multiplications for parameter projections dominate the model’s FLOPs. In this architecture, the inter-layer hybrid design is optimized for extreme efficiency, owing to its simplicity in attention computation, model parallelism and sequence parallelism. SWA layers captures precise local patterns, while linear attention layers efficiently compresses long-range information. The model thus provides linear-time training complexity and constant memory usage during inference, regardless of sequence length, offering substantial efficiency gains for long-context processing. In practice, we implement linear attention using a Gated Linear Attention Module yang2024gated (GLA), where the gating vector $\mathbf{g}_{t}$, derived from low-rank projection parameters, enhances expressivity and recurrent modeling capability:
@@
The 76B model integrates linear attention and SWA in a 1:1 intra-layer parallel hybrid configuration, while standard full-attention layers are interleaved at a 1:6 ratio across layers. This intra-layer hybrid design is optimized for stronger language modeling capabilities, as it enables more intricate attention computation and incorporates additional attention parameters and FLOPs. During parallel hybridization, outputs of both attention branches are normalized to ensure consistent scale and to prevent instability in the early stages of training. In addition, 128 learnable sink tokens (xiao2023efficient, ; dong2024hymba, ) are introduced to mitigate the attention-sink phenomenon in softmax attention and to enhance the flexibility of SWA for local modeling. Specifically, 128 trainable embeddings are prepended to the input; these tokens are attended by all other tokens and also attend to each other without causal masking. This functionality is implemented by customizing FlashAttention (daoflashattention, ) kernels. The model also employs Gated Linear Attention Modules, but in this case, the gating vector is tied directly to the key vector (chou2024metala, ; qinhgrn2, ), i.e.,
$\mathbf{g}_{t}=\mathbf{1}-\mathbf{k}_{t},$ (12)
eliminating the need for extra gating parameters. The FFN modules adopt a sparse Mixture-of-Experts (MoE) design: each MoE layer consists of 16 routed experts (top-1 activated) and one shared expert, so that only about 15% of parameters (12B) are activated per token. To stabilize training and control parameter growth, 7 dense FFN layers are preserved out of 28 total layers (specifically at layers 1, 2, 3, 5, 7, 9, and 11). Similarly, spike coding is applied to the activations of all linear projection layers throughout the model.
@@
Our architectural choices are closely aligned with principles observed in biological brains. i) Linear attention modules exhibit modeling properties analogous to human memory, relying on compressed and continuously updated states (gu2025tradeoffs, ). At each step, they retrieve information only from the current memory, showing Markov-like behavior. From a biological perspective, their stateful temporal recurrence can be viewed as a simplified abstraction of dendritic dynamics with multi-branch morphology. ii) The Mixture-of-Experts (MoE) component embodies the principle of modular sparse activation and functional specialization, reminiscent of the distributed and specialized processing found in neural circuits (o2021and, ). iii) Our spike coding scheme draws inspiration from event-driven and adaptively sparse neuronal activation in biological systems (maass1997networks, ). By combining network-level sparsity (MoE) with neuron-level spiking sparsity, our approach enables on-demand allocation of computation and provides a robust two-scale efficiency strategy. Collectively, these results suggest a promising pathway for designing large-model architectures that are both efficient and biologically plausible (10636118, ).
@@ 3 Training Paradigm
The development of our SpikingBrain models relies on a highly efficient and universal conversion pipeline that is compatible with the open-source model ecosystem. Specifically, we start from a Transformer base model and perform targeted transformations on three key modules—hybrid attention modules, MoE modules, and spike-coding modules—supported by multi-stage training (including CPT and SFT) to adapt the model to the new brain-inspired architectures.
3.1 Attention Map Correspondence
@@
Leveraging this attention-map correspondence, we can initialize the QKV projection parameters of SpikingBrain directly from a pre-trained Transformer checkpoint (kasai2021finetuning, ), and this applies to full attention, linear attention, and SWA. By reusing the learned QK similarity and performing lightweight training on a small amount of data, the attention can be adapted to these local or low-rank cases. Since SWA maintains local modeling precision and linear attention retains global interaction, combining them in a hybrid paradigm provides a closer approximation to the original attention map (dass2023vitality, ; chen2021scatterbrain, ; arora2024simple, ). This enables a smoother transition during conversion and accelerates loss convergence.
Finally, due to the generality of this attention-map formulation, we can construct diverse attention patterns—sparse, local, linear, or hybrid—and transfer them from any pre-trained softmax attention model. This allows for flexible trade-offs between performance and efficiency by adjusting factors such as the proportion of retained full-attention layers or the window size of local attention.
@@
3.2 MoE Upcycling
@@ 3.3.1 Step 1: Adaptive Threshold Spiking
For gradient computation in the optimization phase, since the spiking process is essentially an approximate representation of the original membrane potential, the gradient of the spiking process can be skipped (i.e., using straight-through estimation), which maintains the gradient computation process consistent with conventional training.
@@ 3.3.2 Step 2: Spike Coding
$\mathbf{s}_{\text{INT}}$ $=\sum_{t=1}^{T}\mathbf{s}_{t},\quad\mathbf{y}=V_{\text{th}}(\mathbf{x})\cdot\sum_{t=1}^{T}\mathbf{W}\mathbf{s}_{t},$ (27)
@@
  • Binary Spike Coding {0,1}: This is the most basic event-driven spike coding method. Each time a unit spike is fired (with a value of $1$), it represents an activation of the neuron state, and the spike count is accumulated over continuous time steps. This coding scheme is intuitive and simple, with low computational overhead, making it suitable for scenarios with very low spike counts and effectively reducing system complexity. However, when representing large counts, this coding often requires many time steps to complete the representation. Additionally, the lack of directional optimization for spike firing results in a higher firing rate, further limiting system energy efficiency. Meanwhile, this encoding scheme only supports positive integers (unsigned $\mathbf{s}_{\text{INT}}$) which is not applicable to Adaptive Threshold Spiking.
@@
  • Bitwise Spike Coding: Bitwise coding can be viewed as an event sequence encoding method, where integer count values are expanded bit by bit into spike events over time steps. Each time step corresponds to one binary bit of the count value, significantly compressing the time dimension overhead. As shown in Figure 4 (c), this mechanism supports three implementation forms to accommodate different symbolic representations and precision requirements: i) Pure bitwise encoding, suitable for positive integers, provides extremely high temporal compression in high-count scenarios. For instance, a count of 256 requires 256 consecutive time steps in binary coding, 128 in ternary coding, but only 8 steps in 8-bit bitwise encoding. ii) Bidirectional bitwise encoding uses $\pm 1$ to represent each bit, replacing negative part with 1-1. This halves the time step emission rate and reduces the required number of time steps by one (e.g., 7 steps for a count of 256), while maintaining equivalence in representation. iii) Two’s complement encoding incorporates sign information into the highest bit, supporting both positive and negative counts while retaining binary simplicity and biological plausibility. Compared to ternary coding, which still requires many time steps for higher bit counts, the bitwise encoding scheme significantly compresses time steps, reducing total spike emissions by up to 8× (for an 8-bit count). This effectively reduces overall spike communication overhead and computational load while maintaining precision, making it particularly suitable for high-precision, low-power, and time-sensitive neuromorphic computing tasks. Taking Bidirectional bitwise encoding as an example, its encoding formula and the equivalent computation of matrix multiplication can be expressed as:
    $\mathbf{s}_{\text{INT}}$ $=\sum_{t=1}^{T}2^{t-1}\mathbf{s}_{t},\quad\mathbf{y}=V_{\text{th}}(\mathbf{x})\cdot\sum_{t=1}^{T}2^{t-1}\mathbf{W}\mathbf{s}_{t},$ (28)
@@ 3.3.3 Hardware Adaptation and Deployment Potential
3.4 Multi-stage Conversion Pipeline
The continual pre-training process comprises three stages which progressively extend the context window while adapting the model to new architectures. In the first stage, our models are trained on 100B tokens with a sequence length of 8k, aiming to transfer attention patterns toward local or low-rank variants and ensure loss convergence. Subsequently, the second and third stages extend the sequence length to 32k and 128k, respectively, each trained with 20B to 30B tokens. The entire conversion process consumes about 150B tokens in total. Compared to the $\sim$10T tokens typically required for training from scratch, the CPT approach requires only about 2% of the data, enabling efficient adaptation under resource and budget constraints. All three stages use the Matrix (zhang2024mapneo, ) dataset, with long-context data generated through a simple per-domain packing strategy. The RoPE base remains consistent with the base model at 1M.
@@ 4.2 Operator Adaptation
  • JIT Compilation Optimization: During Triton’s just-in-time compilation stage, the linear attention operators (yang2024gated, ) used in SpikingBrain are reorganized at the code level. By refining instruction pipelining and register allocation, the operators achieve a balance between memory access latency and computational density. This stage directly leverages Triton’s hardware-agnostic general optimization strategies, enabling dynamic and efficient execution of compute kernels.
@@ 4.3 Parallel Topology
The memory demands of training large language models often exceed the capacity of a single GPU. To make training such models feasible, it is essential to employ efficient and scalable distributed training techniques alongside other memory reduction methods. These approaches distribute computational and storage loads while maintaining training efficiency. In this section, we present the four parallelization strategies and corresponding distributed training topologies employed to train SpikingBrain models on the MetaX cluster.
@@
Pipeline parallelism divides the model into different stages by layers. Each GPU is assigned a subset of layers, and intermediate activations are communicated between GPUs during the forward and backward passes using peer-to-peer (p2p) communication. We employ the 1F1B (narayanan2021efficientllmmg, ; harlap2018pipedream, ) scheduling algorithm for efficient pipeline execution. Furthermore, incorporating Mixture-of-Experts (MoE) layers interspersed with dense FFN layers in SpikingBrain-76B helps mitigate the memory imbalance issue, where the first pipeline stage typically exhibits higher memory usage compared to other stages.
@@
In Mixture-of-Experts models, most parameters reside within the expert networks. Expert parallelism (rajbhandari2022deepspeed, ) addresses this by partitioning experts across multiple GPUs. This requires all-to-all communication to dispatch and combine tokens. During the training of SpikingBrain-76B, we employ Grouped GEMM kernels to accelerate computation across experts, and an auxiliary loss is used to promote balanced token routing across experts.